OpenStudy (anonymous):

The area of a circle is directly proportional to the square of its radius. If the radius is halved, by what factor is the area multiplied? a) 1/8 b) 2 c) 1/2 d) 4 e) 1/4 Can you deeply and clearly explain to me, please? Thanks.

OpenStudy (anonymous):

If the area of a circle is directly proportional to the square of the radius, then: \[A=k\times r^2\] i.e. A = some constant times \(r^2\). Imagine then, that instead of r, you have \(\frac{r}{2}\). How does that equality change?

OpenStudy (anonymous):

Give that the area of the circle, A is directly proportional to the square of its radius, R. \[A \alpha R ^{2}\] \[A= k R ^{2}\] , where k is some constant. WHen your radius is halved, R/2. SUbstitute R/2 into the original R, \[ k \left( \frac{R}{2} \right) ^{2}\] \[= k \left( \frac{R^{2}}{4} \right) \] \[= \frac{1}{4}k R ^{2}\] Since A = k R^2, Substituting into the above equation, we will get 1/4 A. Area of the circle decreases by a factor of 1/4